Phil collects .50 stamps and .60 stamps. At the post office he purchased 27 stamps worth $15.00. HOw many of each kind did he get?
Answer by asha(30)

Phil collects .50 and .60 stamps. he purchased 27 stamps for $15.
Let him buy x stamps of .50 and y stamps of .60.
x + y= 27 ---- eqn.1
.50x +.60y =15 --- eqn.2
multiplying eqn.1 by 5 we get
5x +5y = 135--- eqn.3
multiplying eqn.2 by 10 we get
5x+6y =150 --- eqn.4
subtracting eqn.3 from eqn.4we get
5x-5x+6y-5y=150-135
y=15
plugging value of y in eqn.1 we get
x+ 15=27
x=27-15=12
Phil collects 12 stamps of .50 and 15 stamps of .60
Good Luck !!